- Status Report
- August 13, 2022
ISS On-Orbit Status 5 Dec 2002
All ISS systems continue to function nominally, except as noted previously
After wake-up at 4:00am EST (back-shifted from yesterday by another hour),
the crew made preparations for more medical assessments, starting with
a test of the CMRS (crew medical restraint system) by CDR Ken Bowersox
and FE-2/SO Don Pettit. [This device can be secured to the ISS structure
within two minutes, in order to provide a patient restraint surface for
performing emergency medical procedures, such as during ACLS (advanced
cardiac life support). It can also be used to transport a patient between
the station and the Orbiter middeck. It isolates the crew and equipment
electrically during defibrillations and pacing electrical discharges,
accommodates the patient in the supine zero-G positions, provides cervical
spine stabilization and can also restrain two CMOs (crew medical officers)
during their delivery of medical care.]
Throughout the day, all crewmembers completed their first general PFE
(physical fitness evaluation), which checks up on blood pressure and electrocardiogram
during programmed exercise on the CEVIS in the Lab. CDR Bowersox, a qualified
CMO (crew medical officer), assisted with the examination. Readings were
taken with the BP/ECG (blood pressure/electrocardiograph). [BP/ECG provides
automated noninvasive systolic and diastolic blood pressure measurements
while also monitoring and displaying accurate heart rates on a continual
basis at rest and during exercise. Sox plugged it in the defibrillator
outlet in the Lab, since it requires USOS voltage for operation, whereas
the CHeCS HRM (Crew Health Maintenance System heart rate monitor), which
measures and records heart rate only, can operate on 24V for use on the
exercise equipment in the Russian segment, such as the TVIS treadmill.]
The first round of the renal (kidney) stone experiment program continued.
Don Pettit, who started his turn yesterday, today began his urine sample
collections, several times during the day. FE-1 Nikolai Budarin had his
first day, reviewing the CBT (computer-based training) disk and starting
his diet log. Bowersox will join in tomorrow.
Budarin completed the periodic routine inspection of the BRPK water condensate
separator in the Service Module (SM). Later, he also performed the routine
tasks of SOSh life support systems maintenance and IMS inventory update
Ken Bowersox reconfigured the Lab CCAA (common cabin air assembly) air
conditioner for the reduced crew size.
Later, Bowersox and Budarin installed the remaining Metox (metal oxide)
canisters for regeneration and conducted a dry-run of the EMU spacesuit
systems, in preparation for next week’s spacewalk (12/12). [Due to the
failed EV1 biomed harness during 11A, which had a damaged "pigtail"
connector, and the concern that there may be similar electrical cables
in the EMU, today’s spacesuit dry-run was conducted in unpressurized condition,
with no oxygen flowing, until a safety review can be completed on the
ground. All suit donning steps were performed except helmet installation,
with the O2 actuator remaining OFF and no biomed instrumentation used.
This allowed verification of suit fit, i.e., making sure that heel-to-shoulder
and crotch-to-shoulder dimensions provide light contact and the feet remain
seated in the bottom of the boot. Boot sizing inserts (BSIs) may be more
difficult to don on-orbit and may be a little snug.]
All crewmembers again had an hour scheduled for ISS familiarization and
acclimatization, as will be the daily rule for the first two weeks after
During yesterday’s reconfiguration of the SSC (station support computer)
with new hard drives, the LAN (local area network) experienced a momentary
"crash". It was quickly reestablished in an impromptu recovery
by Bowersox, who reloaded it with great attention to detail, and the ground
took notes to update procedures for future use.
Troubleshooting is underway by the ground on the HRF (human research
facility) RIC (rack interface controller) software, after Don Pettit ran
into a problem with it yesterday during reload. Don was requested to provide
a download of log files of the HRF laptop via OCA path.
In order to collect some SSRMS/Canadarm2 thermal trend data, the ground
powered up the Lab RWS (robotics workstation) and SSRMS at 11:45am EST.
To support this activity, the crew was asked to connect the Lab RWS DCP
(display and control panel) bypass power cable beforehand.
MCC-Moscow was planning to conduct today a remote test on both subsets
of the FGB Kurs-P nadir system for automated rendezvous and docking.
Landing of STS-113/Endeavour at KSC/Florida was waved off for the second
day in a row, due to inclement weather at the Cape. Mission managers plan
to try again tomorrow (Friday 12/6), with first landing opportunity at
1:57pm EST. Edwards AFB in California is not being called up for tomorrow,
so Endeavour is either to land in Florida tomorrow or wait until Saturday.